# Instruction-tuning Aligns LLMs to the Human Brain

### v1

@@ Abstract
Instruction-tuning is a widely-adopted method of finetuning that enables large language models (LLMs) to generate output that more closely resembles human responses to natural language queries, in many cases leading to human-level performance on diverse testbeds. However, it remains unclear whether instruction-tuning truly makes LLMs more similar to how humans process language. We investigate the effect of instruction-tuning on LLM-human similarity in two ways: (1) brain alignment, the similarity of LLM internal representations to neural activity in the human language system, and (2) behavioral alignment, the similarity of LLM and human behavior on a reading task. We assess 25 vanilla and instruction-tuned LLMs across three datasets involving humans reading naturalistic stories and sentences. We discover that instruction-tuning generally enhances brain alignment by an average of $6\%$, but does not have a similar effect on behavioral alignment. To identify the factors underlying LLM-brain alignment, we compute correlations between the brain alignment of LLMs and various model properties, such as model size, various problem-solving abilities, and performance on tasks requiring world knowledge spanning various domains. Notably, we find a strong positive correlation between brain alignment and model size (r = 0.95), as well as performance on tasks requiring world knowledge (r = 0.81). Our results demonstrate that instruction-tuning LLMs improves both world knowledge representations and brain alignment, suggesting that mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain.
@@ 1 Introduction
Instruction-tuning is a widely adopted method for finetuning large language models (LLMs) on datasets containing task-specific instructions. This approach enhances their ability to generalize effectively to previously unseen tasks by learning to follow provided instructions (Wang et al., 2022c). Instruction-tuning often costs only a small fraction of compute relative to pretraining (Chung et al., 2022), yet propels pretrained LLMs to incredible performance leaps on reasoning and problem-solving benchmarks. This transformation has enabled LLMs to approach human performance on many tasks, despite using only few (or zero) training examples, as well as to tackle open-world reasoning tasks previously only achievable by humans (Zhang et al., 2023).
@@
In addition to teaching LLMs to understand and follow human instructions, instruction-tuning also improves the ability of LLMs to mimic the ground-truth outputs (often human-written) of the training data. This property allows them to produce more controllable and predictable output that is deemed (1) more desirable by human evaluators on various metrics (Zhang et al., 2023; Chung et al., 2022; Wang et al., 2022b), (2) more aligned to human values (Chia et al., 2023), and (3) more stylistically similar to human outputs (Dasgupta et al., 2022; Safdari et al., 2023).
@@
Consequently, instruction-tuning yields LLMs more similar to humans in both capability and output rememblance. From a neuroscience perspective, this begs the question: Does instruction-tuning make LLMs more similar to the human language system? Previous work has shown that models with higher task performance are more aligned to the human language system (Schrimpf et al., 2021; Goldstein et al., 2022; Caucheteux & King, 2022), and, on some datasets, even hit the estimated noise ceiling. [^1] In fMRI recordings, an upper limit of representational similarity can be computed by sampling from the same participant twice, deducing a threshold defined by the noise level of the data gathering process. However, there has been no similar study on how instruction-tuning, the method that has enabled powerful LLMs such as ChatGPT, affects alignment to the human language system.
@@
In this work, we explore the impact of instruction-tuning on the alignment between LLMs and the human language system, considering two aspects: (1) brain alignment, which assesses how closely LLMs’ internal representations match neural activity patterns in the human language system, and (2) behavioral alignment, which evaluates the similarity between LLM behavior and human behavior. To conduct this study, both LLMs and human participants are presented with the same language stimuli comprised of naturalistic stories and sentences. For LLMs, we analyze their internal representations and per-word perplexity, while for humans, we use previously collected brain activity data from functional magnetic resonance imaging (fMRI) experiments and per-word reading times.
@@
To measure brain alignment, we use the Brain-Score (Schrimpf et al., 2018) linear predictivity metric, assessing how well LLM representations predict human brain activity in response to the same language stimuli (Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021; Oota et al., 2023), using data from three neural datasets: Pereira et al. (2018), Blank et al. (2014), and Wehbe et al. (2014). To evaluate behavioral alignment, we use a benchmark in Brain-Score which calculates the Pearson correlation between LLM per-word perplexity and human per-word reading times from the Futrell et al. (2018) dataset. Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs and humans share similarities in terms of which words and sentences they find challenging or surprising. Because models vary in their brain and behavioral alignment across different architectures and training objectives (Schrimpf et al., 2021), we estimate the effect of instruction-tuning across 17 instruction-tuned LLMs and 8 vanilla LLMs, and report a significant increase in brain alignment by instruction-tuned models compared to vanilla ones.
@@
To investigate why instruction-tuning increases alignment to human brain activity, we then estimate the contribution of various LLM properties towards brain alignment. Specifically, we compute Pearson correlations between an LLM’s brain alignment and its properties, including next-word prediction (NWP) ability, model size, a range of problem-solving abilities, and world knowledge spanning different domains. The evaluation of the latter two properties is based on the Big-Bench Hard benchmark (BBH) (Suzgun et al., 2022) and the Massive Multi-task Language Understanding benchmark (MMLU) (Hendrycks et al., 2021), respectively.
@@
    Instruction-tuning generally improves the alignment of LLM representations to brain activity, increasing brain alignment by 6.26.2% on average for the LLMs and neural datasets we tested (Figure 1).
@@
    Investigating the factors underlying LLM-brain alignment, we find that world knowledge and model size are strongly correlated with brain alignment (r = $0.81$ and $0.95$ for instruction-tuned models, respectively; Figure 2).
@@
    Surprisingly, our results generally indicate that instruction-tuning LLMs does not enhance behavioral alignment with human reading times. Furthermore, behavioral alignment on this dataset demonstrates poor correlations with all other measures we investigate, including task performance and model size (Figure 3).
Figure 1: Instruction-tuning aligns LLM representations to human brain activity. (A) We evaluate brain alignment as the similarity of an LLM’s internal representations to human brain activity, using a linear predictivity metric in Brain-Score. We evaluate 25 vanilla and instruction-tuned LLMs with sizes between 77M and 33B parameters. We use 3 neural datasets of humans reading naturalistic stories and sentences: Pereira2018, Blank2014, and Wehbe2014, and compute the average across the three datasets. (B) Instruction-tuning improves average brain alignment by 6.2% on average. Each point above the identity line represents an instruction-tuned LLM that has greater brain alignment than its vanilla counterpart. Error bars (here and elsewhere) represent median absolute deviation over human participants. (C) Instruction-tuning generally improves brain alignment on all three neural datasets. (D) We instruction-tune LLaMA-7B using the Alpaca dataset (“Instruction” model). We also train an ablation model with the same process and training data, but remove the instruction portion from each training sample (“No Instruction” model). This experiment shows that brain alignment improvements are due to both (1) training data (present in both models) and (2) the process of training LLMs to understand and follow instructions (present only in original model).
@@ 2 Background & Related Work
LM properties linked to Brain alignment.   There is a growing body of work disentangling the contribution of various LM properties towards brain alignment. These include studying how brain alignment is driven by next-word prediction ability (Schrimpf et al., 2021; Caucheteux & King, 2022), multi-word semantics (Merlin & Toneva, 2022), performance on various NLP tasks (Oota et al., 2022), and model size (Antonello et al., 2023). To disentangle the contribution of various LM properties toward brain alignment, we use instruction-tuned LLMs. They are especially useful as they have been trained to respond to a standard question-answer format, allowing us to evaluate LLMs on a wide array of tasks and in a more fine-grained manner. We expand this area of research by identifying that world knowledge is a key property underlying LLM-brain alignment.
@@ 3 Language Models
We evaluate the brain alignment of 25 large language models (LLMs) from two model families: T5 (Raffel et al., 2020) and LLaMa (Touvron et al., 2023). T5 models are encoder-decoder LLMs pre-trained on the Colossal Common Crawl Corpus (C4), a corpus of 356 billion tokens, using a masked infilling objective, and then further finetuned on multi-task mixture of unsupervised and supervised tasks converted into a text-to-text format. In our study, we use all five T5 models with sizes between 77M to 11B parameters. LLaMA models (Touvron et al., 2023) are decoder-only LLMs trained on 1.6 trillion tokens from a mixture of corpora including C4, English CommonCrawl, Wikipedia, Github, and more. For LLaMA, we use the 7B, 13B, and 33B parameter versions in our study.
For the instruction-tuned variants of T5 models, we utilize a variety of models finetuned on the FLAN suite (1515M examples for 1,836 different tasks accompanied by instructions, Chung et al., 2022), Alpaca (5252K instruction-following examples generated through methods inspired by Self-Instruct, Wang et al. (2022a), Taori et al., 2023), and GPT4ALL (437437K instruction-following examples generated with GPT-3.5-turbo, Anand et al., 2023) datasets. As for the LLaMa model family, we employ Vicuna’s 77B, 1313B, and 3333B models (Chiang et al., 2023), which are finetuned on user-shared conversations. Additionally, we incorporate the StableVicuna-13B model, which further refines the Vicuna-13B model using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) on a range of conversational and instructional datasets. We also use the 77B version of Alpaca (Taori et al., 2023). Additional details about these LLMs can be found in Appendix A.
Table 1: Example of instruction-tuning training data format: (Instruction, Input, Output) from the Alpaca dataset (Taori et al., 2023). The input field is optional for certain types of instructions.
@@
“Write a short paragraph “The importance of “The use of renewable energy is growing
about the given topic.” using renewable energy.” rapidly in relevance and …”
@@ 4 Brain Alignment
Brain alignment refers to the method of evaluating the similarity between LLM representations and human brain activity (Figure 1). This relies on fMRI recordings of human subjects while they read language stimuli on potentially any topic (here: Pereira et al., 2018; Blank et al., 2014; Wehbe et al., 2014). The same language stimuli from prior brain recordings are provided as input to LLMs, whose intermediate layer activations are extracted as their representations of the language stimuli. We follow a general approach previously used in several works (Schrimpf et al., 2018; 2021; Jain & Huth, 2018; Toneva & Wehbe, 2019; Oota et al., 2023; Aw & Toneva, 2023). Specifically, we use the linear predictivity metric implemented in Brain-Score (Schrimpf et al., 2020, Figure 1), first training a linear function to predict fMRI voxels associated with the human language system using LLM representations as input features. We then apply this linear function to held-out brain activity data from the original corpus of recordings, and evaluate the brain alignment of the LLM as the Pearson correlation between the predicted and actual brain activity data. For each LLM, we evaluate its brain alignment for every layer (e.g., LLaMA-7B has 32 layers), and use the highest value as the LLM’s brain alignment value, following Schrimpf et al. (2018).
Datasets   We use three fMRI datasets to measure the brain alignment of LLMs. Each dataset involves a different set of human participants and uses a different set of language stimuli.
Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, nine participants read 384 sentences organized into 96 text passages. In experiment 3, six participants read 243 sentences in 72 text passages. Each sentence was displayed for four seconds on a screen.
Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorceror’s Stone (Rowling et al., 1998). Participants read the chapter at a fixed interval of one word every 0.5 seconds.
@@ 4.1 Instruction-tuning aligns LLM representations to human brain activity
First, we study the effect of instruction-tuning on LLM brain alignment. We compute each LLM’s average brain alignment as the mean of its brain alignment on the 3 neural datasets. We find that instruction-tuning improves alignment by an average of 6.2% across all tested LLMs (Figure 1B). This holds across all three neural datasets, with average improvements of +6.9% on Pereira2018, +8.0% improvement on Blank2014, and +3.8% on Wehbe2014 (Figure 1C). This provides a strong signal that instruction-tuning improves brain alignment. Moreover, a smaller instruction-tuned model can attain higher brain alignment than a larger model from the same family that has not been instruction-tuned (e.g., Alpaca-7B v.s. LLaMa-13B, see detailed results in Appendix E).
@@
To longitudinally study how instruction-tuning aligns LLM representations to brain activity, we separately instruction-tune a LLaMA-7B model on the Stanford Alpaca instruction dataset (Taori et al., 2023) for 3 epochs. By evaluating checkpoints regularly during training, we find that instruction-tuning progressively improves brain alignment (Figure 1D). We also perform an ablation study to disambiguate the effect on brain alignment of (1) the instruction-following ability provided by the instruction-tuning step and (2) added training data. We fine-tune LLaMA-7B with the same process and data, but remove the instruction portion from each training sample. We observe that brain alignment of this ablated model increases during fine-tuning but stays lower than its instruction-following counterpart (Figure 1D). This shows that brain alignment improvements are due to both factors.
@@
Figure 2: World knowledge and model size are important factors underlying LLM-brain alignment. To identify factors underlying brain alignment, we test Pearson correlations between brain alignment and various LLM properties, such as model size, world knowledge in various domains (MMLU benchmark), and various types of problem-solving abilities (BBH benchmark). Insets display results on individual datasets, with stars reflecting statistical significance (n.s. = p >> 0.05, * = p << 0.05, ** = p << 0.005, etc.) (A) Brain alignment is significantly and strongly correlated with world knowledge as evaluated by the MMLU Overall score (r = 0.81), which reports the mean performance across all MMLU subjects. (B) Brain alignment is significantly and strongly correlated with performance on the world knowledge task category in BBH (r = 0.68). (C) Brain alignment is significantly and strongly correlated with model size (logarithm of number of parameters) (r = 0.95).

### v2

@@ Abstract
Instruction-tuning is a widely adopted finetuning method that enables large language models (LLMs) to generate output that more closely resembles human responses. However, no studies have shown that instruction-tuning actually teaches LLMs to process language in a similar manner as humans. We investigate the effect of instruction-tuning on aligning LLM and human language processing mechanisms in two ways: (1) brain alignment, the similarity of LLM internal representations to neural activity in the human language system, and (2) behavioral alignment, the similarity of LLM and human behavior on a reading task. We assess 25 vanilla and instruction-tuned LLMs on three datasets involving humans reading naturalistic stories and sentences, and find that instruction-tuning generally enhances brain alignment ($\sim$6%), but has no similar effect on behavioral alignment. To identify factors underlying this improvement in brain alignment, we compute correlations between brain alignment and various LLM properties, such as model size, problem-solving, and world knowledge understanding. Notably, we find a strong positive correlation between brain alignment and model size (r = 0.95), as well as performance on tasks requiring world knowledge (r = 0.81). Our results demonstrate that instruction-tuning LLMs improves both world knowledge representations and brain alignment, suggesting that the mechanisms that encode world knowledge in LLMs also improve representational alignment to the human brain.
@@ 1 Introduction
Instruction-tuning is a widely adopted method that finetunes large language models (LLMs) on datasets containing many examples of different tasks and their descriptive instructions, enhancing their ability to generalize to previously unseen tasks by learning to follow provided instructions (Wang et al., 2022c). Despite costing only a small fraction of compute relative to pretraining (Chung et al., 2022), instruction-tuning yields large performance improvements on reasoning and problem-solving benchmarks. This generalization has allowed LLMs to tackle open-world reasoning tasks previously achievable only by humans (Zhang et al., 2023) while only using a few (or zero) task-specific training examples.
@@
In addition to teaching LLMs to understand and follow human instructions, instruction-tuning also improves the ability of LLMs to mimic human-written ground-truth outputs. This fluency allows them to produce more controllable and predictable output that is deemed (1) more desirable to human evaluators (Zhang et al., 2023; Chung et al., 2022; Wang et al., 2022b), (2) more aligned to human values (Chia et al., 2023), and (3) more stylistically similar to human outputs (Dasgupta et al., 2022; Safdari et al., 2023).
@@
Consequently, instruction-tuning yields LLMs more similar to humans in both capability and output resemblance. From a neuroscience perspective, this begs the question: Does instruction-tuning make LLMs more functionally similar to the human language system? Previous work has shown that models with higher task performance are more aligned to the human language system (Schrimpf et al., 2021; Goldstein et al., 2022; Caucheteux & King, 2022), hitting the estimated noise ceiling [^1] In fMRI recordings, a noise ceiling for representational similarity can be computed by sampling from the same participant twice, obtaining an upper limit for how well an ideal model could perform, defined by the internal consistency of neural responses and noise level of the data gathering process. on some datasets. However, there has been no similar study on how instruction-tuning affects alignment with the human language system.
@@
In this work, we measure the effect of instruction-tuning on the alignment between language mechanisms in LLMs and the human brain in two ways: (1) brain alignment, an “internal” metric that assesses how well internal feature representations of LLMs match neural representations in the human language system, and (2) behavioral alignment, an “external” metric which evaluates the similarity between LLM and human behavioral measurements. In our study, both LLMs and humans are presented with the same language stimuli comprised of naturalistic stories and sentences. For LLMs, we record their internal representations and per-word perplexity. For humans, we use previously recorded brain activity data from functional magnetic resonance imaging (fMRI) experiments and per-word reading times.
@@
To measure brain alignment, we use the Brain-Score (Schrimpf et al., 2018) linear predictivity metric, assessing how well LLM representations predict human brain activity in response to the same language stimuli (Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021; Oota et al., 2023), using data from three neural datasets: Pereira et al. (2018), Blank et al. (2014), and Wehbe et al. (2014). To evaluate behavioral alignment, we use a benchmark in Brain-Score which calculates the Pearson correlation between LLM per-word perplexity and human per-word reading times from the Futrell et al. (2018) dataset. Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs match humans in their patterns of words and sentences they find challenging or surprising. Because models vary in their brain and behavioral alignment across different architectures and training objectives (Schrimpf et al., 2021), we estimate the effect of instruction-tuning by evaluating 8 vanilla LLMs and 17 LLMs that were further instruction-tuned, and report a significant increase in brain alignment due to instruction-tuning.
@@
To investigate why instruction-tuning increases alignment to human brain activity, we study the relationships between brain alignment and various LLM properties. Specifically, we compute Pearson correlations between an LLM’s brain alignment and its properties, including next-word prediction (NWP) ability, model size, a range of problem-solving abilities, and world knowledge spanning different domains. We evaluated the latter two properties using the Big-Bench Hard benchmark (BBH; Suzgun et al., 2022) and the Massive Multi-task Language Understanding benchmark (MMLU; Hendrycks et al., 2021), respectively.
@@
    Instruction-tuning generally aligns LLM representations to human brain activity, increasing brain alignment by $6.2$% on average (Figure 1).
@@
    Investigating the factors underlying LLM-brain alignment, we find that brain alignment is strongly correlated with world knowledge (r = $0.81$) and model size (r = $0.95$) (Figure 2).
@@
    Surprisingly, our results indicate that instruction-tuning LLMs generally does not enhance behavioral alignment with human reading times. Furthermore, behavioral alignment on this dataset is poorly correlated with all other measures we investigate, including task performance and model size (Figure 3).
Figure 1: Instruction-tuning aligns LLM representations to human brain activity. (A) The same language stimuli are presented to LLMs and human participants. Next, we fit a linear regression model from LLM layer activations to fMRI responses in the human language system. We apply this linear model to predict held-out fMRI responses from the original corpus of recordings, and compute brain alignment as the Pearson correlation between the predicted and actual fMRI responses. We evaluate 25 vanilla and instruction-tuned LLMs with sizes between 77M and 33B parameters. We compute the average across 3 neural datasets of humans reading naturalistic stories and sentences: Pereira2018, Blank2014, and Wehbe2014. (B) Instruction-tuning improves average brain alignment by 6.2% on average. Each point above the identity line is an instruction-tuned LLM with greater brain alignment than its vanilla version. Error bars, here and elsewhere, represent median absolute deviation over human participants. (C) Instruction-tuning improves average brain alignment on all three datasets. (D) We instruction-tune LLaMA-7B on the Alpaca dataset (“Instruction” model) and train an ablation model with the same data, but without the instruction in each training sample (“No Instruction” model). Our results show that brain alignment improvements are due to both (1) training data (present in both models) and (2) training LLMs to understand and follow instructions (present only in the first model).
@@ 2 Background & Related Work
LM properties linked to Brain alignment.   One exciting research area focuses on disentangling the contribution of various LM properties towards brain alignment. These include studying how brain alignment is driven by next-word prediction ability (Schrimpf et al., 2021; Caucheteux & King, 2022), multi-word semantics (Merlin & Toneva, 2022), performance on various NLP tasks (Oota et al., 2022), and model size (Antonello et al., 2023). To contribute to this growing body of work, we use instruction-tuned LLMs. They are especially useful as they have been trained to respond to a standard question-answer format, allowing us to evaluate LLMs on a wide array of tasks and in a more fine-grained manner. We expand this area of research by demonstrating that world knowledge is a key property underlying LLM-brain alignment.
@@ 3 Language Models
We evaluate the brain alignment of 25 large language models (LLMs) from two model families: T5 (Raffel et al., 2020) and LLaMa (Touvron et al., 2023a). T5 models are encoder-decoder LLMs pre-trained using a masked infilling objective on the Colossal Common Crawl Corpus (C4), a corpus of 356 billion tokens, and then further finetuned on a multi-task mixture of unsupervised and supervised tasks converted into a text-to-text format. We use all five T5 models, with sizes between 77M to 11B. LLaMA models (Touvron et al., 2023a) are decoder-only LLMs trained on 1.6 trillion tokens from a mixture of corpora including C4, English CommonCrawl, Wikipedia, GitHub. We use the 7B, 13B, and 33B versions.
For the instruction-tuned variants of T5 models, we use a variety of models finetuned on FLAN ($15$M examples for 1,836 different tasks accompanied by instructions, Chung et al., 2022), Alpaca ($52$K instruction-following examples generated with methods inspired by Self-Instruct, Wang et al. (2022a), Taori et al., 2023), and GPT4ALL ($437$K instruction-following examples generated with GPT-3.5-turbo, Anand et al., 2023) datasets. For the instruction-tuned variants of LLaMa, we use Vicuna’s $7$B, $13$B, and $33$B models (Chiang et al., 2023), which are finetuned on user-shared conversations. We also use StableVicuna-13B, which further refines Vicuna-13B using reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) on a range of conversational and instructional datasets. We also use the $7$B version of Alpaca (Taori et al., 2023). We provide more details in Appendix A.
We also tested 8 models from the GPT2 family; four vanilla models (Small, Medium, Large, XL) and their corresponding versions further instruction-tuned on Alpaca. However, the models performed close to random on reasoning and world knowledge benchmarks (BBH, MMLU). We hypothesize this is due to their size (all $<$ 1.5B parameters), making fine-tuning on a relatively small instruction-following dataset insufficient. Moreover, GPT2 was not originally pretrained to perform instruction-following tasks (unlike T5 models). We include their MMLU, BBH, brain and behavioral alignment results in Appendix G, H, and L.
@@
“Write a short paragraph “The importance of “The use of renewable energy
about the given topic.” using renewable energy.” is growing…” (paragraph)
Table 1: Example of Instruction-tuning training data format: (Instruction, Input, Output) from the Alpaca dataset (Taori et al., 2023). The input field is optional for certain instructions.
@@ 4 Brain Alignment
Brain alignment refers to the method of evaluating the similarity between LLM representations and human brain activity (Figure 1A and Appendix C). This relies on fMRI recordings of human subjects while they read language stimuli on potentially any topic (here: Pereira et al., 2018; Blank et al., 2014; Wehbe et al., 2014). The same language stimuli from prior brain recordings are provided as input to LLMs, whose intermediate layer activations are extracted as their representations of the language stimuli. We follow a general approach previously used in several works (Schrimpf et al., 2018; Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021; Oota et al., 2023; Aw & Toneva, 2023). Specifically, we use the linear predictivity metric implemented in Brain-Score (Schrimpf et al., 2018), which fits a linear regression model from LLM layer activations to fMRI responses in the human language system. We then apply this linear model to predict held-out fMRI responses from the original corpus of recordings, and compute brain alignment as the Pearson correlation between the predicted and actual fMRI responses. For each LLM, we evaluate its brain alignment for every layer (e.g., LLaMA-7B has 32 layers), and use the highest value as the LLM’s brain alignment value, following Schrimpf et al. (2018).
Datasets   We use three fMRI datasets to measure the brain alignment of LLMs. Each dataset involves a different set of human participants and a different set of language stimuli.
Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, 9 participants read 384 sentences taken from 96 text passages. In experiment 3, 6 participants read 243 sentences from 72 text passages. Each sentence was displayed for 4 seconds on a screen.
Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to 8 naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorcerer’s Stone (Rowling et al., 1998). Participants read the chapter at a fixed interval of one word every 0.5 seconds.
@@ 4.1 Instruction-tuning aligns LLM representations to human brain activity
First, we study the effect of instruction-tuning on LLM brain alignment. We compute each LLM’s average brain alignment as the mean of its brain alignment on the 3 neural datasets. We find that instruction-tuning improves brain alignment by an average of 6.2% across tested LLMs (Figure 1B). This holds across all three neural datasets, with improvements of +6.9% on Pereira2018, +8.0% improvement on Blank2014, and +3.8% on Wehbe2014 (Figure 1C). Moreover, a smaller instruction-tuned model can attain higher brain alignment than a larger vanilla model from the same family (e.g., Alpaca-7B v.s. LLaMa-13B, see Appendix G). However, some of the larger models (T5-XL, LLaMA) have minimal or no increases, possibly because the vanilla models are already close to the noise ceiling (Appendix G).
@@
Next, to longitudinally study how instruction-tuning aligns LLM representations to brain activity, we separately instruction-tune a LLaMA-7B model on the Stanford Alpaca instruction dataset (Taori et al., 2023) for 3 epochs. By evaluating checkpoints regularly during training, we find that instruction-tuning progressively improves brain alignment (Figure 1D, Appendix K). We also disambiguate the effect on brain alignment of (1) the instruction-following ability gained from instruction-tuning and (2) added training data. We fine-tune LLaMA-7B with the same process and data, but remove the instruction from each training sample. Brain alignment of this ablated model increases during fine-tuning but stays lower than its instruction-following counterpart (Figure 1D), indicating that brain alignment improvements are due to both factors. In Appendix J, we also report preliminary results using recent models from the Gemma (Team et al., 2024) and LLaMA-2 (Touvron et al., 2023b) families and observe similar trends. Further, these trends hold across different similarity metrics besides Linear predictivity, e.g., RSA and CKA (Appendix J).
@@
Figure 2: World knowledge and model size are important factors underlying LLM-brain alignment. To identify factors underlying brain alignment, we test Pearson correlations between brain alignment and various LLM properties, such as model size, world knowledge in various domains (MMLU benchmark), and various types of problem-solving abilities (BBH benchmark). In the figure, insets display results on individual datasets, with stars reflecting statistical significance (n.s. = p $>$ 0.05, * = p $<$ 0.05, ** = p $<$ 0.005, etc.) (A) Brain alignment is significantly and strongly correlated with world knowledge as evaluated by the MMLU Overall score (r = 0.81), which reports the mean performance across all MMLU subjects. (B) Brain alignment is significantly and strongly correlated with performance on the world knowledge task category in BBH (r = 0.68). (C) Brain alignment is significantly and strongly correlated with model size (logarithm of number of parameters) (r = 0.95). In Appendix I, we provide a larger version of this figure with labels for each data point.
Benchmark, Task category Correlation (r) to Adjusted No. Avg. Model
Brain Alignment p-value tasks Accuracy
MMLU, Overall Score 0.809 $<$ 0.0005 57 0.36
MMLU, STEM 0.792 $<$ 0.0005 18 0.28
MMLU, Humanities 0.791 $<$ 0.0005 13 0.34
MMLU, Social Sciences 0.807 $<$ 0.0005 12 0.41
MMLU, Others 0.809 $<$ 0.0005 14 0.40
BBH, Overall score 0.384 0.18 23 0.28
BBH, Algorithmic reasoning 0.194 0.56 8 0.22
BBH, Language understanding 0.163 0.59 3 0.43
BBH, World knowledge 0.679 $<$ 0.005 5 0.36
BBH, Multilingual reasoning -0.035 0.90 1 0.19
BBH, Others 0.478 0.08 6 0.27
Table 2: Brain alignment strongly correlates with world knowledge in all subject domains in MMLU and the world knowledge category in BBH. Brain alignment is not significantly correlated with other problem-solving abilities in BBH (e.g., algorithmic or multilingual reasoning). We obtain p-values after performing false discovery rate (FDR) correction.